OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x – y = –4 4x – 3y = –6

OpenStudy (anonymous):

2x – y = –4 4x – 3y = –6 Those are our equations. We can solve for y on the first equation: 2x – y = –4 We add y both sides--->2x=-4+y We add 4 both sides--->2x+4=y y=2x+4 How do we continue now?